# Brain Foundation Models: A Survey on Advancements in Neural Signal Processing and Brain Discovery

### v1

@@ Conclusion
In this survey, we introduce the concept of BFMs as a transformative paradigm in neural signal processing and brain discovery. By providing the first formal definition and theoretical framework for BFMs, we delineate how they differ from conventional foundation models, emphasizing their specialized design for handling complex, noisy, and heterogeneous neural data. We synthesize the latest architectural innovations and offer novel perspectives on their applications. Our analysis highlights that BFMs not only extend the frontiers of brain decoding by generalizing robustly across diverse scenarios and tasks but also enable advanced brain simulation and discovery through digital twin brain construction. This dual capability underscores their potential to revolutionize both clinical diagnostics and neuroscientific research. Furthermore, we outline a forward-looking roadmap identifying critical challenges and future directions. These include improving data integration and quality, developing novel training strategies and multimodal architectures, enhancing model interpretability, and addressing ethical, regulatory, and standardization concerns. Overcoming these challenges is essential to fully harness the transformative potential of BFMs in personalized medicine, neurorehabilitation, and beyond. Ultimately, this survey serves as a foundational reference for researchers, practitioners, and policymakers. By bridging the gap between neuroscience and artificial intelligence, BFMs are poised to reshape our understanding of brain function and drive the development of next-generation brain technologies.
@@
Xinliang Zhou received the B.E. degree in Mechatronics Engineering from Beijing Jiaotong University, Beijing, China, in 2021, and the Ph.D. degree in Computer Science from Nanyang Technological University, Singapore, in 2025. His research interests include Brain-Computer Interfaces, Brain Foundation Models, and Interpretable Artificial Intelligence.
@@
Chenyu Liu received the B.E. degree in Software Engineering from the University of Electronic Science and Technology of China, Sichuan, China, in 2020. He is currently working toward the Ph.D. degree with the College of Computing and Data Science, Nanyang Technological University, Singapore. His research interests include Brain-Computer Interface, Multivariate Time Series, and Multimodal Spatio-Temporal Data Mining.
@@
Zhisheng Chen is currently working toward a B.E. degree in Remote Sensing Science and Technology at China University of Mining and Technology. His research interests include Brain-Computer Interfaces and Large-scale Neural Signal Modeling.
@@
Kun Wang received the Ph.D. degree in Computer Science from University of Science and Technology of China (USTC), Anhui, China, in 2024. He has published papers in top conferences and journals, such as ICLR and IEEE Transactions on Knowledge and Data Engineering. His primary research interests include ML4Science, Foundation Models and AI Agents.
Yi Ding earned a Ph.D. in Computer Science and Engineering from Nanyang Technological University, Singapore, in 2023, a master’s degree in Electrical and Electronics Engineering from the same university in 2018, and a bachelor’s degree in Information Science and Technology from Donghua University, Shanghai, China, in 2017. He is a Research Assistant Professor at the College of Computing and Data Science, Nanyang Technological University. His research interests encompass Brain-Computer Interfaces, Deep Learning, Neural Signal Processing, Foundation Models, and Affective Computing.
@@
Ziyu Jia received the Ph.D. degree in Computer Science and Technology from Beijing Jiaotong University, Beijing, China, in 2022. He is an Assistant professor at the Brainnetome Center, Institute of Automation, Chinese Academy of Sciences. His interests include Novel Time Series Data Mining, Machine Learning, and Deep Learning Algorithms. He is also interested in various physiological time series applications in affective Brain-Computer Interfaces and Sleep Stage Classification.
Qingsong Wen is the Head of AI & Chief Scientist at Squirrel Ai Learning. Before that, he worked at Alibaba, Qualcomm, Marvell, etc., and received his M.S. and Ph.D. degrees in Electrical and Computer Engineering from Georgia Institute of Technology, USA. His research interests include Machine Learning, Data Mining, and Signal Processing, especially AI for Time Series (AI4TS), AI for Education (AI4EDU), Decision Intelligence, LLM and GenAI. He has published over 100 top-ranked AI conference and journal papers, had multiple Oral/Spotlight Papers at NeurIPS, ICML, and ICLR, had multiple Most Influential Papers at IJCAI, received multiple IAAI Deployed Application Awards at AAAI, and won First Place in the SP Grand Challenge at ICASSP. Currently, he serves as Organizer/Co-Chair of Workshop on AI for Time Series, Workshop on AI for Education, and Workshop on AI Agent for Information Retrieval, at KDD, AAAI, IJCAI, CIKM, ICDM, etc. In addition, he regularly serves as Area Chair of the top conferences, including NeurIPS, KDD, ICASSP, etc. He also serves as Associate Editor for Neurocomputing, Associate Editor for IEEE Signal Processing Letters, Guest Editor for Applied Energy, and Guest Editor for IEEE Internet of Things Journal. He is a committee member of the IEEE Machine Learning for Signal Processing Technical Committee (MLSP TC).

### v2

@@ Conclusion
In this survey, we introduce the concept of BFMs as a transformative paradigm in neural signal processing and brain discovery. By providing the first formal definition and theoretical framework for BFMs, we delineate how they differ from conventional foundation models, emphasizing their specialized design for handling complex, noisy, and heterogeneous neural data. We synthesize the latest architectural innovations and offer novel perspectives on their applications. Our analysis highlights that BFMs not only extend the frontiers of brain decoding by generalizing robustly across diverse scenarios and tasks but also enable advanced brain simulation and discovery through digital twin brain construction. This dual capability underscores their potential to revolutionize both clinical diagnostics and neuroscientific research. Furthermore, we outline a forward-looking roadmap identifying critical challenges and future directions. These include improving data integration and quality, developing novel training strategies and multimodal architectures, and enhancing model interpretability. Overcoming these challenges is essential to fully harness the transformative potential of BFMs in personalized medicine, neurorehabilitation, and beyond. Ultimately, this survey serves as a foundational reference for researchers, practitioners, and policymakers. By bridging the gap between neuroscience and artificial intelligence, BFMs are poised to reshape our understanding of brain function and drive the development of next-generation brain technologies.
@@
Xinliang Zhou received the B.E. degree in Mechatronics Engineering from Beijing Jiaotong University, Beijing, China, in 2021, and the Ph.D. degree in Computer Science from Nanyang Technological University, Singapore, in 2025. His research interests include Brain-Computer Interfaces, Brain Foundation Models, and Interpretable Artificial Intelligence. (Email Address: xinliang001@e.ntu.edu.sg)
@@
Chenyu Liu received the B.E. degree in Software Engineering from the University of Electronic Science and Technology of China, Sichuan, China, in 2020. He is currently working toward the Ph.D. degree with the College of Computing and Data Science, Nanyang Technological University, Singapore. His research interests include Brain-Computer Interface, Multivariate Time Series, and Multimodal Spatio-Temporal Data Mining. (Email Address: chenyu003@e.ntu.edu.sg)
@@
Zhisheng Chen received the B.E. degree in Remote Sensing Science and Technology from China University of Mining and Technology. He is currently working toward the Master’s degree with the Institute of Computing Technology, Chinese Academy of Sciences. His research interests include Brain-Computer Interfaces, Large-scale Neural Signal Modeling, and Brain Foundation Models. (Email Address: zhishengchen@cumt.edu.cn)
@@
Kun Wang received the Ph.D. degree in Computer Science from University of Science and Technology of China (USTC), Anhui, China, in 2024. He is a Research Fellow at the College of Computing and Data Science, Nanyang Technological University. His primary research interests include ML4Science, Foundation Models and AI Agents. (Email Address: wang.kun@ntu.edu.sg)
@@
Ziyu Jia received the Ph.D. degree in Computer Science and Technology from Beijing Jiaotong University, Beijing, China, in 2022. He is an Assistant professor at the Brainnetome Center, Institute of Automation, Chinese Academy of Sciences. His interests include Novel Time Series Data Mining, Machine Learning, and Deep Learning Algorithms. He is also interested in various physiological time series applications in affective Brain-Computer Interfaces and Sleep Stage Classification. (Email Address: jia.ziyu@outlook.com)
Yi Ding earned a Ph.D. in Computer Science and Engineering from Nanyang Technological University, Singapore, in 2023, a master’s degree in Electrical and Electronics Engineering from the same university in 2018, and a bachelor’s degree in Information Science and Technology from Donghua University, Shanghai, China, in 2017. He is a Research Assistant Professor at the College of Computing and Data Science, Nanyang Technological University. His research interests encompass Brain-Computer Interfaces, Deep Learning, Neural Signal Processing, Foundation Models, and Affective Computing. (Email Address: ding.yi@ntu.edu.sg)
Qingsong Wen is Head of AI & Chief Scientist at Squirrel AI Learning. He previously held positions at Alibaba, Qualcomm, and Marvell, and earned his M.S. and Ph.D. in Electrical and Computer Engineering from the Georgia Institute of Technology. His research focuses on Machine Learning, Data Mining, and Signal Processing, particularly in AI for Time Series (AI4TS), AI for Education (AI4EDU), LLMs, Generative AI, and AI Agents. He serves as Area Chair for conferences including NeurIPS and KDD, Associate Editor for IEEE TPAMI, IEEE SPL, and Neurocomputing. He is a member of the IEEE Machine Learning for Signal Processing Technical Committee (MLSP TC). (Email Address: qingsongedu@gmail.com)